# Spatial-Sign based Maxsum Test for High Dimensional Location Parameters

### v1

@@ 4 Simulation
Here we consider the scatter matrix ${\bf\Sigma}=(0.5^{|i-j|})_{1\leq i,j\leq p}$. Two sample sizes $n=50,100$ and three dimensions $p=200,400,600$ are considered. All the findings in this section are derived from 1000 repetitions. Table 1 presents the empirical sizes of the six tests mentioned above. It was observed that the spatial-sign based tests–SS-MAX, SS-SUM, and SS-CC–are able to effectively manage the empirical sizes in a majority of scenarios. However, when dealing with distributions that are not multivariate normal, the SUM test tends to have empirical sizes that fall below the nominal level. Similarly, the COM test also exhibits smaller sizes when operating under non-normal distributions.
@@ 7.1 Proof of main lemmas
where $F$ represent the distribution of $\bm{X}$, $\bm{\eta}=(\eta_{1},\cdots,\eta_{2p})$. Similar with Hettmansperger and Randles (2002), the linearisation of this equation shows
@@
Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then, if $\log p=o(n^{1/3})$,
@@ 7.1.1 Proof of Lemma 1(Bahadur representation)
Similar with Cheng et al. (2023), by Assumption 2 and Markov inequality, we have that: $\max R_{i}^{-2}=O_{p}(\zeta_{1}^{2}n^{1/2})$, $\max\delta_{1i}=O_{p}(\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|^{2}\max\hat{R}_{i}^{-2}=O_{p}(n^{-1/2})$ and $\max\delta_{2i}=O_{p}(n^{-1/4})$.
@@
[…]
@@ 7.2.4 Proof of Theorem 5(Without elliptical distribution)
By Lemma 6 and Equation 7.10, $\mathbb{E}((\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j})^{2})\leq O(n^{-1}\log p\text{ tr}(\mathbf{\Sigma}_{u}^{2}))=o(p^{-2}\text{tr}(\mathbf{R}^{2}))$. Then we obtain $G_{n1}=o_{p}(\sigma_{n})$. Similar to $A_{n1}$, we can show $G_{n2}=o_{p}(\sigma_{n})$. Taking the same procedure as $A_{n2}$, we can obtain $A_{n3}=o_{p}(\sigma_{n})$. Similarly to the processing of Equation 7.1, we get
@@ 7.2.5 Proof of Theorem 6(Asymptotically independent under $H_{0}$)
Suppose $\{\bm{Y}_{1},\bm{Y}_{2},\cdots,\bm{Y}_{n}\}$ are sample from $N(0,\mathbb{E}\bm{U}_{1}^{\top}\bm{U}_{1})$ , and independent with $\bm{U}_{1},\cdots,\bm{U}_{n}$ or write as $\bm{\xi}_{1},\cdots,\bm{\xi}_{n})$. The key idea is to show that: $(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}),V(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}))$ has the same limiting distribution as $W(\bm{Y}_{1},\cdots,\bm{Y}_{n}),V(\bm{Y}_{1},\cdots,\bm{Y}_{n}))$.
@@
$+\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right),$
for some $\delta\in(0,1)$. Again, due to $\mathbb{E}\left(\bm{\xi}_{t}\right)=\mathbb{E}\left(\bm{Y}_{t}\right)=0$ and $\mathbb{E}\left(\bm{\xi}_{t}\bm{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\bm{Y}_{t}\bm{Y}_{t}^{\top}\right)$, we can verify that
@@
$\left|\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\right|\leq C\beta^{2},$
@@
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\delta\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right)\right|$
@@
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d+1,0}+\delta\left(\bm{z}_{d+1}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d+1,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d+1,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d+1,v}-\bm{z}_{d,0,v}\right)\right|$
@@
holds with probability approaching one. Consequently, we have that: with probability one,
@@
Similar with Equation 7.23, for Equation 7.19-(v),
@@
$\sum_{d=1}^{n}\left|\mathrm{E}\left\{f\left(W_{d},V_{d}\right)\right\}-\mathrm{E}\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\right|\leq C\beta^{2}n^{-1/2}\log^{3}(n)+C^{\prime}n^{-1/2}\rightarrow 0,$
@@ 7.2.6 Proof of Theorem 7(asymptotically independent under $H_{1,local}$)
Similar with the proof in Theorem 6, it’s suffice to show the result holds for normal version, i.e. it suffice to show that:
@@ 7.3 Some useful lemmas
(Lemma A4. in Cheng et al. (2023))Suppose Assumptions 1-3 holds with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then,

### v2

@@ 4 Simulation
Here we consider the scatter matrix ${\bf\Sigma}=(0.5^{|i-j|})_{1\leq i,j\leq p}$. Two sample sizes $n=50,100$ and three dimensions $p=200,400,600$ are considered. All the findings in this section are derived from 1000 repetitions. Table 1 presents the empirical sizes of the six tests mentioned above. It was observed that the spatial-sign based tests–SS-MAX, SS-SUM, and SS-CC–are able to effectively manage the empirical sizes in a majority of scenarios. Under the normality assumption, the Type I error of the MAX method increases as the ratio $p/n$ grows. This may be due to the component $n\bar{x}_{i}/\hat{\sigma}_{ii}$ of the max statistic following a $t(n)$ distribution, which deviates from the normal distribution. In contrast, the SS-MAX method is more suitable for heavy-tailed data. When $p$ is fixed and $n$ increases, leading to a closer approximation to the normal distribution, the MAX method exhibits improved control over the Type I error. However, when dealing with distributions that are not multivariate normal, the SUM test tends to have empirical sizes that fall below the nominal level. Similarly, the COM test also exhibits smaller sizes when operating under non-normal distributions.
@@ 5.1 US stocks data
5.2 Paired colon dataset
Another important application of the one-sample test discussed in this paper is assessing the mean difference between two paired samples. In this section, we utilize our methods to test the mean difference for paired samples and consider the colon dataset provided by Alon et al. (1999). The colon dataset includes gene expression data from 40 colon cancer patients, comprising 22 paired samples from normal and tumor colon tissue and additional 18 samples from tumor tissue, with each sample containing 2,000 gene expressions. Our objective is to assess whether the mean gene expression levels differ between normal and tumor tissues. To streamline the analysis, we exclude unpaired samples and retain only the $n=22$ paired normal and tumor tissue samples. To compare the methods, MAX, SUM and COM methods are also displayed. We observed that SUM test fails to reject the null hypothesis while others strongly reject if the significant level is set to $\alpha=0.05$, see Figure 4. It is aligned with the simulation results of non-normal cases, suggesting a significant difference in gene expression between normal and tumor tissues, warranting further investigation. In addition, we find that the COM test successfully rejects the null hypothesis, whereas the SUM test does not. This indicates that methods based on the theorem of the independence between the test statistic can enhance test power while ensuring that Type I error remains controlled, particularly when data sparsity is uncertain.
Table 4: The p-values of testing the difference of gene expression levels of the normal and tumor colon tissues.
SS-MAX SS-SUM SS-CC MAX SUM COM
$6.88\times 10^{-5}$ $6.71\times 10^{-7}$ $1.33\times 10^{-6}$ $1.68\times 10^{-3}$ $9.62\times 10^{-2}$ $3.30\times 10^{-3}$
@@ 6 Conclusion
Acknowledgement
The authors thank the editor, the associate editor and three anonymous referees for helpful comments and discussions. This paper is partially supported by Shenzhen Wukong Investment Company, Tianjin Science Fund for Outstanding Young Scholar (23JCJQJC00150), the Fundamental Research Funds for the Central Universities under Grant No. ZB22000105 and 63233075, the China National Key R&D Program (Grant Nos. 2019YFC1908502, 2022YFA1003703, 2022YFA1003802, 2022YFA1003803) and the National Natural Science Foundation of China Grants (Nos. 12271271, 11925106, 12231011, 11931001 and 11971247). Long Feng and Ping Zhao are the co-corresponding authors.
@@ 7.1 Proof of main lemmas
where $F$ represent the distribution of $\bm{X}$, $\bm{\eta}=(\eta_{1},\cdots,\eta_{2p})$. Similar to Hettmansperger and Randles (2002), the linearisation of this equation shows
@@
Suppose Assumptions 1-3 hold with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then, if $\log p=o(n^{1/3})$,
@@ 7.1.1 Proof of Lemma 1(Bahadur representation)
Similar to Cheng et al. (2023), by Assumption 2 and Markov inequality, we have that: $\max R_{i}^{-2}=O_{p}(\zeta_{1}^{2}n^{1/2})$, $\max\delta_{1i}=O_{p}(\|\hat{\mathbf{D}}^{-1/2}\bm{\theta}\|^{2}\max\hat{R}_{i}^{-2}=O_{p}(n^{-1/2})$ and $\max\delta_{2i}=O_{p}(n^{-1/4})$.
@@
$\frac{1}{n}\zeta_{1}^{-1}\sum_{i=1}^{n}\hat{R}_{i}^{-1}$
$=$ $\frac{1}{n}\zeta_{1}^{-1}\sum_{i=1}^{n}R_{i}^{-1}(1+O_{p}(n^{-1/2}(\log p)^{1/2}))$
$=$ $\left\{1+O_{p}(n^{-1/2})\right\}\left\{1+O_{p}(n^{-1/2}(\log p)^{1/2})\right\}$
$=$ $1+O_{p}(n^{-1/2}(\log p)^{1/2}).$
@@ 7.2.4 Proof of Theorem 5(Without elliptical distribution)
By Lemma 6 and Equation 7.10, $\mathbb{E}((\bm{U}_{i}^{\top}\left(\hat{\mathbf{D}}_{ij}^{-1/2}\mathbf{D}^{1/2}-\mathbf{I}_{p}\right)\bm{U}_{j})^{2})\leq O(n^{-1}\log p\text{ tr}(\mathbf{\Sigma}_{u}^{2}))=o(p^{-2}\text{tr}(\mathbf{R}^{2}))$. Then we obtain $G_{n1}=o_{p}(\sigma_{n})$. Similar to $A_{n1}$, we can show $G_{n2}=o_{p}(\sigma_{n})$. Taking the same procedure as $A_{n2}$, we can obtain $A_{n3}=o_{p}(\sigma_{n})$. Similar to the processing of Equation 7.1, we get
@@ 7.2.5 Proof of Theorem 6(Asymptotically independent under $H_{0}$)
Suppose $\{\bm{Y}_{1},\bm{Y}_{2},\cdots,\bm{Y}_{n}\}$ are sample from $N(0,\mathbb{E}\bm{U}_{1}^{\top}\bm{U}_{1})$ , and independent with $\bm{U}_{1},\cdots,\bm{U}_{n}$(or write as $\bm{\xi}_{1},\cdots,\bm{\xi}_{n})$. The key idea is to show that: $(W(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}),V(\bm{\xi}_{1},\cdots,\bm{\xi}_{n}))$ has the same limiting distribution as $(W(\bm{Y}_{1},\cdots,\bm{Y}_{n}),V(\bm{Y}_{1},\cdots,\bm{Y}_{n}))$.
@@
$+\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\widetilde{\delta}\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right),$
for some $\widetilde{\delta}\in(0,1)$. Again, due to $\mathbb{E}\left(\bm{\xi}_{t}\right)=\mathbb{E}\left(\bm{Y}_{t}\right)=0$ and $\mathbb{E}\left(\bm{\xi}_{t}\bm{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\bm{Y}_{t}\bm{Y}_{t}^{\top}\right)$, we can verify that
@@
$\left|\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\widetilde{\delta}\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\right|\leq C\beta^{2},$
@@
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d,0}+\widetilde{\delta}\left(\bm{z}_{d}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d,v}-\bm{z}_{d,0,v}\right)\right|$
@@
$\left|\frac{1}{6}\sum_{l=1}^{n}\sum_{k=1}^{n}\sum_{v=1}^{n}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\bm{z}_{d+1,0}+\widetilde{\delta}\left(\bm{z}_{d+1}-\bm{z}_{d,0}\right)\right)\left(\bm{z}_{d+1,l}-\bm{z}_{d,0,l}\right)\left(\bm{z}_{d+1,k}-\bm{z}_{d,0,k}\right)\left(\bm{z}_{d+1,v}-\bm{z}_{d,0,v}\right)\right|$
@@
hold with probability approaching one. Consequently, we have that: with probability one,
@@
Similar to Equation 7.23, for Equation 7.19-(v),
@@
$\sum_{d=1}^{n}\left|\mathrm{E}\left\{f\left(W_{d},V_{d}\right)\right\}-\mathrm{E}\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\right|\leq C\beta^{2}n^{-1/2}\log^{3}n+C^{\prime}n^{-1/2}\rightarrow 0,$
@@ 7.2.6 Proof of Theorem 7(asymptotically independent under $H_{1,local}$)
Similar to the proof in Theorem 6, it’s suffice to show the result holds for normal version, i.e. it suffice to show that:
@@ 7.3 Some useful lemmas
(Lemma A4. in Cheng et al. (2023))Suppose Assumptions 1-3 hold with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$. Then,